OpenStudy (anonymous):

how would i find the rate of change of the diagonals in a rectangle? when i tried it i got D= sqrt(l^2+w^2) then i took the dd/dt and got (.5(l^2+w^2)^-.5)(2l(dl/dt)+2w(dw/dt))

OpenStudy (amistre64):

\[D= \sqrt{(l^2+w^2)}\] \[D'= \frac{(l^2+w^2)'}{2\sqrt{(l^2+w^2)}}\] \[D'= \frac{(l^2)'+(w^2)'}{2\sqrt{(l^2+w^2)}}\] \[D'= \frac{2l+2w}{2\sqrt{(l^2+w^2)}}\] \[D'= \frac{2(l+w)}{2\sqrt{(l^2+w^2)}}\] \[D'= \frac{l+w}{2\sqrt{(l^2+w^2)}}\] maybe

OpenStudy (amistre64):

.... i forgot to delete the 2 in the denom. brains said delete and fingers said ... nah keep it

OpenStudy (anonymous):

i dont really understand the first step

OpenStudy (amistre64):

the first step, and i left a few things out as well just ... is simply deriving the sqrt function itself

OpenStudy (amistre64):

and as a result; the derivate of its innards pops out

OpenStudy (anonymous):

oh ok ty

OpenStudy (amistre64):

let me type it up better, since i really just flubbed it, from here \[D'= \frac{(l^2)'+(w^2)'}{2\sqrt{(l^2+w^2)}}\] \[D'= \frac{2l\ l'+2w\ w'}{2\sqrt{(l^2+w^2)}}\] \[D'= \frac{l\ l'+w\ w'}{\sqrt{(l^2+w^2)}}\] is a better rendition, the 's being the derivatives or rates of change of the length and width